OpenStudy (anonymous):

Stable High School has a total of 112 boys and girls who play sports. If the number of boys, b is 16 more than twice the number of girls, g, how many boys play sports at this high school? Which system of equations would help you solve this problem? b + g = 112 b = 2g + 16 b + g = 112 g = 2b + 16 b − g = 112 b = 2g − 16 None of the systems above would solve this problem.

OpenStudy (ajprincess):

b+g=112 b=2g+16

OpenStudy (callisto):

@ajprincess you should explain your answer...

OpenStudy (ajprincess):

Sure @callisto nd thanx for remindng t.

OpenStudy (ajprincess):

Total means the sum of both boys and girls That is b+g=112 Number of boys is 16 more than twice of the number of girls. That is number of boys=2*no f girls+16 b=2g+16